# Random Feature Representation Boosting

### v1

@@ Acknowledgements
TC has been supported by the EPSRC Programme Grant EP/S026347/1. NZ has been supported by the Roth Scholarship at Imperial College London. We acknowledge computational resources and support provided by the Imperial College Research Computing Service (DOI: 10.14469/hpc/2232). For the purpose of open access, the author(s) has applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Analytic Solutions to Sandwiched Least Squares
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},\Delta\in\mathbb{R},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(\Delta)=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}\Delta X_{i}\big{\|}^{2}+\lambda\Delta^{2}$
@@
$\Delta_{\textnormal{scalar}}$ $=\frac{\langle R,XW\rangle_{F}}{\|XW\|_{F}^{2}+n\lambda}=\frac{\frac{1}{n}\sum_{i=1}^{n}\langle W^{\top}X_{i},R_{i}\rangle}{\frac{1}{n}\sum_{i=1}^{n}\|W^{\top}X_{i}\|^{2}+\lambda}.$
@@
We first rewrite the objective $J(\Delta)$ using the Frobenius norm
$J(\Delta)=\frac{1}{n}\|X\Delta W-R\|_{F}^{2}+\lambda\Delta^{2}.$
Differentiating with respect to $\Delta\in\mathbb{R}$ and setting $J^{\prime}(\Delta)=0$ gives
$0=J^{\prime}(\Delta)=\frac{2}{n}(XW)^{\top}(\Delta XW-R)+2\lambda\Delta$
$\iff\Delta\|XW\|^{2}_{F}+n\lambda\Delta=\langle XW,R\rangle_{F},$
from which the result follows by factoring out $\Delta$. Note that the second derivative is
$J^{\prime\prime}(\Delta)=\frac{2}{n}\|XW\|^{2}_{F}+2\lambda>0,$
@@
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},\Delta=\textnormal{diag}(\delta_{1},...,\delta_{D})\in\mathbb{R}^{D\times D},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(\Delta)=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}\Delta X_{i}\big{\|}^{2}+\lambda\|\Delta\|^{2}_{F}$
@@
$A=WW^{\top}\odot X^{\top}X,\qquad\qquad b=\textnormal{diag}(WR^{\top}X).$
@@
We expand $J(\Delta)$ and find that
$J(\Delta)$ $=J(\delta_{1},...,\delta_{D})$
$=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}\Delta X_{i}\big{\|}^{2}+\lambda\|\Delta\|^{2}_{F}$
$=\frac{1}{n}\sum_{i=1}^{n}\big{(}R_{i}-W^{\top}[\delta_{1}X_{i,1},...,\delta_{D}X_{i,D}]^{\top}\big{)}^{\top}\big{(}R_{i}-W^{\top}[\delta_{1}X_{i,1},...,\delta_{D}X_{i,D}]^{\top}\big{)}+\lambda\sum_{k=1}^{D}\delta_{k}^{2}.$
Differentiating with respect to a specific $\delta_{k}$ gives
00 0 $=\frac{\partial J(\delta_{1},...,\delta_{D})}{\partial\delta_{k}}=\frac{1}{n}\sum_{i=1}^{n}-2\big{(}W_{k}X_{i,k}\big{)}^{\top}\big{(}R_{i}-W^{\top}[\delta_{1}X_{i,1},...,\delta_{D}X_{i,D}]^{\top}\big{)}+2\lambda\delta_{k}$
$\qquad\iff\sum_{i=1}^{n}(W_{k}X_{i,k})^{\top}R_{i}=\sum_{i=1}^{n}\bigg{[}(W_{k}X_{i,k})^{\top}W^{\top}\Delta X_{i}+\lambda\delta_{k}\bigg{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{i=1}^{n}\bigg{[}\lambda\delta_{k}+\sum_{j=1}^{D}(W_{k}X_{i,k})^{\top}W_{j}X_{i,j}\delta_{j}\bigg{]}.$
This implies that the value of $\Delta$ which minimizes the objective is the solution to the system of linear equations given by
$b=(A+\lambda I)\Delta,$
@@
$A=WW^{\top}\odot X^{\top}X,\qquad\qquad b=\textnormal{diag}(WR^{\top}X).$
This solution is unique since the objective is strictly convex with hessian given by $\frac{2}{n}A+2\lambda I$. ∎
@@
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},\Delta\in\mathbb{R}^{D\times p},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(\Delta)$ $=\frac{1}{n}\sum_{i=1}^{n}\big{\|}r_{i}-W^{\top}\Delta x_{i}\big{\|}^{2}+\sum_{k=1}^{D}\sum_{j=1}^{p}\lambda\Delta_{k,j}^{2}$
$=\frac{1}{n}\|W^{\top}\Delta X^{\top}-R^{\top}\|^{2}_{F}+\lambda\|\Delta\|^{2}_{F}$
@@
$\Delta_{\textnormal{dense}}=U\bigg{[}U^{\top}WR^{\top}XV\oslash\bigg{(}\lambda n\mathbf{1}+\textnormal{diag}(\Lambda^{W})\otimes\textnormal{diag}(\Lambda^{X})\bigg{)}\bigg{]}V^{\top},$
@@
00 0 $=\nabla J(\Delta)$
$=\frac{-2}{n}W(W^{\top}\Delta X^{\top}-R^{\top})X+2\lambda\Delta$
$\iff$ $WR^{\top}X=WW\Delta X^{\top}X+\lambda N\Delta.$
Letting $WW^{\top}=U\Lambda^{W}U^{\top}$ and $X^{\top}X=V\Lambda^{X}V^{\top}$ be spectral decompositions, and setting $\widetilde{\Delta}=U^{\top}\Delta V$ we see that
$U^{\top}WR^{\top}XV=\Lambda^{W}\widetilde{\Delta}\Lambda^{X}+\lambda n\widetilde{\Delta}.$
@@
$(U^{\top}WR^{\top}XV)_{k,j}=\Lambda^{W}_{k,k}\widetilde{\Delta}_{k,j}\Lambda^{X}_{j,j}+\lambda n\widetilde{\Delta}_{k,j},$
@@
$\widetilde{\Delta}_{k,j}=\frac{(U^{\top}WR^{\top}XV)_{k,j}}{\Lambda^{W}_{k,k}\Lambda^{X}_{j,j}+N\lambda},$
from which the conclusion follows after change of basis $\Delta=U\widetilde{\Delta}V^{\top}$. The solution is unique since the problem is strictly convex, which is easily proved using the convexity of the Frobenius norm and the linearity of the matrix expressions involving $\Delta$, together with the strict convexity of the regularization term. ∎
@@
3Delta = (U.T @ W @ R.T @ X @ V)
4Delta = Delta / (n*lambda_reg + SW[:, None]*SX[None, :])
5Delta = U @ Delta @ V.T
@@ Appendix B Functional Gradient Inner Product
$\operatorname*{argmin}\limits_{\Delta\in\mathbb{R}^{D\times p}\textnormal{ such that }\|\Delta f\|_{L_{2}^{D}(\mu_{n})}\leq 1}\langle h,\Delta f\rangle_{L^{2}(\mu_{n})}$
@@
$\frac{1}{n}\sum_{i=1}^{n}\|h(x_{i})-\Delta f(x_{i})\|^{2},\qquad\textnormal{subject to}\qquad\frac{1}{n}\sum_{i=1}^{n}\|\Delta f(x_{i})\|^{2}=1.$
@@
$\langle h,\Delta f\rangle_{L_{2}^{D}(\mu)}=\frac{1}{n}\sum_{i=1}^{n}\langle h(x_{i}),\Delta f(x_{i})\rangle=\frac{1}{n}\sum_{i=1}^{n}\langle H_{i},\Delta F_{i}\rangle.$ (7)
@@
$\|\Delta f\|^{2}_{L_{2}^{D}(\mu_{n})}=\frac{1}{n}\sum_{i=1}^{n}\|\Delta f(x_{i})\|^{2}=\frac{1}{n}\sum_{i=1}^{n}\|\Delta F_{i}\|^{2}=1,$ (8)
with equality instead of inequality, since we always obtain a bigger inner product in magnitude by normalizing by $\|\Delta f\|^{2}_{L_{2}^{D}(\mu_{n})}$. Minimizing (7) subject to (8) is equivalent to solving a constrained least squares problem, since we can write
$\frac{1}{n}\sum_{i=1}^{n}\|H_{i}-\Delta F_{i}\|^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\|H_{i}\|^{2}-2\langle H_{i},\Delta F_{i}\rangle+\|\Delta F_{i}\|^{2},$
where we see that the first term $\|H_{i}\|^{2}$ is constant w.r.t $\Delta$, and the third term gives the constraint. Hence, minimizing the constrained least squares problem is equivalent to maximizing the inner product, and the solution to the original problem is obtained by multiplying the least squares solution by $-1$ since we are interested in the argmin rather than the argmax.
@@
$J(\Delta,\nu)$ $=\frac{1}{n}\sum_{i=1}^{n}|H_{i}-\Delta F_{i}|^{2}-\nu\left(\frac{1}{n}\sum_{i=1}^{n}|\Delta F_{i}|^{2}-1\right)$
$=\frac{1}{n}\|H^{\top}-\Delta F^{\top}\|^{2}_{F}-\nu\left(\frac{1}{n}\|\Delta F^{\top}\|^{2}_{F}-1\right).$
Differentiating with respect to $\Delta$ gives
$0=\nabla_{1}J(\Delta,\nu)=\frac{-2}{n}(H^{\top}-\Delta F^{\top})F-\nu\frac{2}{n}\Delta F^{\top}F,$
@@
If $\nu=1$, then $H^{\top}F=0$, implying that $\langle H_{i},\Delta F_{i}\rangle_{L_{2}^{D}(\mu)}=0$ for all matrices $\Delta$. Hence the same closed-form solution holds in this case too. ∎
@@
It is clear from the proof how to augment the expression for $\Delta$ when $F$ is not of full rank. However, in practice we instead use ridge regression for increased numerical stability. A similar result as the above can be proven for ridge regression, albeit with a more complicated expression for $\Delta$ involving non-trivial combinations of $\Lambda$ and $\lambda$. We omit this detail here, and simply use ridge regression in practice.
@@ C.3 Categorical Cross-Entropy Loss
$l(x,y)=-\log(p_{y}(x))$
where $p$ is the softmax function defined by
$p_{y}(x)=\frac{\exp(x_{y})}{\sum_{j=1}^{K}\exp(x_{j})}.$
@@
$\frac{\partial l(x,y)}{\partial x_{k}}$ $=-\frac{\partial}{\partial x_{k}}\log(p_{y}(x))$
$=-\frac{1}{p_{y}(x)}\frac{\partial}{\partial x_{k}}p_{y}(x)$
$=-\frac{1}{p_{y}(x)}\frac{\partial}{\partial x_{k}}\frac{\exp(x_{y})}{\sum_{j=1}^{K}\exp(x_{j})}$
$=-\frac{1}{p_{y}(x)}\left(\frac{\mathbf{1}_{y=k}\exp(x_{y})\sum_{j=1}^{K}\exp(x_{j})-\exp(x_{y})\exp(x_{k})}{\left(\sum_{j=1}^{K}\exp(x_{j})\right)^{2}}\right)$
$=-\frac{1}{p_{y}(x)}\left(\mathbf{1}_{y=k}p_{y}(x)-p_{y}(x)p_{k}(x)\right)$
$=p_{k}(x)-\mathbf{1}_{y=k}$
@@
where $P(X)\in\mathbb{R}^{n\times K}$ is the matrix given by $P(X)_{i,k}=p_{k}(W^{\top}\Phi(x_{i}))$, and $E_{Y}\in\mathbb{R}^{n\times K}$ is the matrix given by $(E_{Y})_{i,k}=\mathbf{1}_{y_{i}=k}$.

### v2

@@ Acknowledgements
TC has been supported by the EPSRC Programme Grant EP/S026347/1. NZ has been supported by the Roth Scholarship at Imperial College London. We acknowledge computational resources and support provided by the Imperial College Research Computing Service (DOI: 10.14469/hpc/2232). For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Appendix A Analytic Solutions to Sandwiched Least Squares
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},A\in\mathbb{R},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(A)=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}AX_{i}\big{\|}^{2}+\lambda A^{2}$
@@
$A_{\textnormal{scalar}}$ $=\frac{\langle R,XW\rangle_{F}}{\|XW\|_{F}^{2}+n\lambda}=\frac{\frac{1}{n}\sum_{i=1}^{n}\langle W^{\top}X_{i},R_{i}\rangle}{\frac{1}{n}\sum_{i=1}^{n}\|W^{\top}X_{i}\|^{2}+\lambda}.$
@@
We first rewrite the objective $J(A)$ using the Frobenius norm
$J(A)=\frac{1}{n}\|XAW-R\|_{F}^{2}+\lambda A^{2}.$
Differentiating with respect to $A\in\mathbb{R}$ and setting $J^{\prime}(A)=0$ gives
$0=J^{\prime}(A)=\frac{2}{n}(XW)^{\top}(AXW-R)+2\lambda A$
$\iff A\|XW\|^{2}_{F}+n\lambda A=\langle XW,R\rangle_{F},$
from which the result follows by factoring out $A$. Note that the second derivative is
$J^{\prime\prime}(A)=\frac{2}{n}\|XW\|^{2}_{F}+2\lambda>0,$
@@
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},A=\textnormal{diag}(a_{1},...,a_{D})\in\mathbb{R}^{D\times D},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(A)=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}AX_{i}\big{\|}^{2}+\lambda\|A\|^{2}_{F}$
@@
$C=WW^{\top}\odot X^{\top}X,\qquad\qquad b=\textnormal{diag}(WR^{\top}X).$
@@
We expand $J(A)$ and find that
$J(A)$ $=J(a_{1},...,a_{D})$
$=\frac{1}{n}\sum_{i=1}^{n}\big{\|}R_{i}-W^{\top}AX_{i}\big{\|}^{2}+\lambda\|A\|^{2}_{F}$
$=\frac{1}{n}\sum_{i=1}^{n}\big{(}R_{i}-W^{\top}[a_{1}X_{i,1},...,a_{D}X_{i,D}]^{\top}\big{)}^{\top}\big{(}R_{i}-W^{\top}[a_{1}X_{i,1},...,a_{D}X_{i,D}]^{\top}\big{)}+\lambda\sum_{k=1}^{D}a_{k}^{2}.$
Differentiating with respect to a specific $a_{k}$ gives
00 0 $=\frac{\partial J(a_{1},...,a_{D})}{\partial a_{k}}=\frac{1}{n}\sum_{i=1}^{n}-2\big{(}W_{k}X_{i,k}\big{)}^{\top}\big{(}R_{i}-W^{\top}[a_{1}X_{i,1},...,a_{D}X_{i,D}]^{\top}\big{)}+2\lambda a_{k}$
$\qquad\iff\sum_{i=1}^{n}(W_{k}X_{i,k})^{\top}R_{i}=\sum_{i=1}^{n}\bigg{[}(W_{k}X_{i,k})^{\top}W^{\top}AX_{i}+\lambda a_{k}\bigg{]}$
$\qquad\qquad\qquad\qquad\qquad\qquad=\sum_{i=1}^{n}\bigg{[}\lambda a_{k}+\sum_{j=1}^{D}(W_{k}X_{i,k})^{\top}W_{j}X_{i,j}a_{j}\bigg{]}.$
This implies that the value of $A$ which minimizes the objective is the solution to the system of linear equations given by
$b=(C+\lambda I)A,$
@@
$C=WW^{\top}\odot X^{\top}X,\qquad\qquad b=\textnormal{diag}(WR^{\top}X).$
This solution is unique since the objective is strictly convex with hessian given by $\frac{2}{n}C+2\lambda I$. ∎
@@
Let $R\in\mathbb{R}^{n\times d},W\in\mathbb{R}^{D\times d},A\in\mathbb{R}^{D\times p},$ and $X\in\mathbb{R}^{n\times D}$. Let $\lambda>0$. Then the minimum of
$J(A)$ $=\frac{1}{n}\sum_{i=1}^{n}\big{\|}r_{i}-W^{\top}Ax_{i}\big{\|}^{2}+\sum_{k=1}^{D}\sum_{j=1}^{p}\lambda A_{k,j}^{2}$
$=\frac{1}{n}\|W^{\top}AX^{\top}-R^{\top}\|^{2}_{F}+\lambda\|A\|^{2}_{F}$
@@
$A_{\textnormal{dense}}=U\bigg{[}U^{\top}WR^{\top}XV\oslash\bigg{(}\lambda n\mathbf{1}+\textnormal{diag}(\Lambda^{W})\otimes\textnormal{diag}(\Lambda^{X})\bigg{)}\bigg{]}V^{\top},$
@@
00 0 $=\nabla J(A)$
$=\frac{-2}{n}W(W^{\top}AX^{\top}-R^{\top})X+2\lambda A$
$\iff$ $WR^{\top}X=WWAX^{\top}X+\lambda NA.$
Letting $WW^{\top}=U\Lambda^{W}U^{\top}$ and $X^{\top}X=V\Lambda^{X}V^{\top}$ be spectral decompositions, and setting $\widetilde{A}=U^{\top}AV$ we see that
$U^{\top}WR^{\top}XV=\Lambda^{W}\widetilde{A}\Lambda^{X}+\lambda n\widetilde{A}.$
@@
$(U^{\top}WR^{\top}XV)_{k,j}=\Lambda^{W}_{k,k}\widetilde{A}_{k,j}\Lambda^{X}_{j,j}+\lambda n\widetilde{A}_{k,j},$
@@
$\widetilde{A}_{k,j}=\frac{(U^{\top}WR^{\top}XV)_{k,j}}{\Lambda^{W}_{k,k}\Lambda^{X}_{j,j}+N\lambda},$
from which the conclusion follows after change of basis $A=U\widetilde{A}V^{\top}$. The solution is unique since the problem is strictly convex, which is easily proved using the convexity of the Frobenius norm and the linearity of the matrix expressions involving $A$, together with the strict convexity of the regularization term. ∎
@@
3A = (U.T @ W @ R.T @ X @ V)
4A = A / (n*lambda_reg + SW[:, None]*SX[None, :])
5A = U @ A @ V.T
@@ Appendix B Functional Gradient Inner Product
$\operatorname*{argmin}\limits_{A\in\mathbb{R}^{D\times p}\textnormal{ such that }\|Af\|_{L_{2}^{D}(\mu_{n})}\leq 1}\langle h,Af\rangle_{L^{2}(\mu_{n})}$
@@
$\frac{1}{n}\sum_{i=1}^{n}\|h(x_{i})-Af(x_{i})\|^{2},\qquad\textnormal{subject to}\qquad\frac{1}{n}\sum_{i=1}^{n}\|Af(x_{i})\|^{2}=1.$
@@
$\langle h,Af\rangle_{L_{2}^{D}(\mu)}=\frac{1}{n}\sum_{i=1}^{n}\langle h(x_{i}),Af(x_{i})\rangle=\frac{1}{n}\sum_{i=1}^{n}\langle H_{i},AF_{i}\rangle.$ (7)
@@
$\|Af\|^{2}_{L_{2}^{D}(\mu_{n})}=\frac{1}{n}\sum_{i=1}^{n}\|Af(x_{i})\|^{2}=\frac{1}{n}\sum_{i=1}^{n}\|AF_{i}\|^{2}=1,$ (8)
with equality instead of inequality, since we always obtain a bigger inner product in magnitude by normalizing by $\|Af\|^{2}_{L_{2}^{D}(\mu_{n})}$. Minimizing (7) subject to (8) is equivalent to solving a constrained least squares problem, since we can write
$\frac{1}{n}\sum_{i=1}^{n}\|H_{i}-AF_{i}\|^{2}$ $=\frac{1}{n}\sum_{i=1}^{n}\|H_{i}\|^{2}-2\langle H_{i},AF_{i}\rangle+\|AF_{i}\|^{2},$
where we see that the first term $\|H_{i}\|^{2}$ is constant w.r.t $A$, and the third term gives the constraint. Hence, minimizing the constrained least squares problem is equivalent to maximizing the inner product, and the solution to the original problem is obtained by multiplying the least squares solution by $-1$ since we are interested in the argmin rather than the argmax.
@@
$J(A,\nu)$ $=\frac{1}{n}\sum_{i=1}^{n}|H_{i}-AF_{i}|^{2}-\nu\left(\frac{1}{n}\sum_{i=1}^{n}|AF_{i}|^{2}-1\right)$
$=\frac{1}{n}\|H^{\top}-AF^{\top}\|^{2}_{F}-\nu\left(\frac{1}{n}\|AF^{\top}\|^{2}_{F}-1\right).$
Differentiating with respect to $A$ gives
$0=\nabla_{1}J(A,\nu)=\frac{-2}{n}(H^{\top}-AF^{\top})F-\nu\frac{2}{n}AF^{\top}F,$
@@
If $\nu=1$, then $H^{\top}F=0$, implying that $\langle H_{i},AF_{i}\rangle_{L_{2}^{D}(\mu)}=0$ for all matrices $A$. Hence the same closed-form solution holds in this case too. ∎
@@
It is clear from the proof how to augment the expression for $A$ when $F$ is not of full rank. However, in practice we instead use ridge regression for increased numerical stability. A similar result as the above can be proven for ridge regression, albeit with a more complicated expression for $A$ involving non-trivial combinations of $\Lambda$ and $\lambda$. We omit this detail here, and simply use ridge regression in practice. Note also that $F(F^{\top}F)^{-1}$ can be expressed as a pseudo-inverse of $F$, after suitable transpositions of the matrices involved.
@@ C.3 Categorical Cross-Entropy Loss
$l(x,y)=-\log(s_{y}(x))$
where $s$ is the softmax function defined by
$s_{y}(x)=\frac{\exp(x_{y})}{\sum_{j=1}^{K}\exp(x_{j})}.$
@@
$\frac{\partial l(x,y)}{\partial x_{k}}$ $=-\frac{\partial}{\partial x_{k}}\log(s_{y}(x))$
$=-\frac{1}{s_{y}(x)}\frac{\partial}{\partial x_{k}}s_{y}(x)$
$=-\frac{1}{s_{y}(x)}\frac{\partial}{\partial x_{k}}\frac{\exp(x_{y})}{\sum_{j=1}^{K}\exp(x_{j})}$
$=-\frac{1}{s_{y}(x)}\left(\frac{\mathbf{1}_{y=k}\exp(x_{y})\sum_{j=1}^{K}\exp(x_{j})-\exp(x_{y})\exp(x_{k})}{\left(\sum_{j=1}^{K}\exp(x_{j})\right)^{2}}\right)$
$=-\frac{1}{s_{y}(x)}\left(\mathbf{1}_{y=k}s_{y}(x)-s_{y}(x)s_{k}(x)\right)$
$=s_{k}(x)-\mathbf{1}_{y=k}$
@@
where $P(X)\in\mathbb{R}^{n\times K}$ is the matrix given by $P(X)_{i,k}=s_{k}(W^{\top}\Phi(x_{i}))$, and $E_{Y}\in\mathbb{R}^{n\times K}$ is the matrix given by $(E_{Y})_{i,k}=\mathbf{1}_{y_{i}=k}$.